A 200kg block is pushed along a horizontal frictionless table a distance of 3meter,by a horizontal force of 12newton. Find
i)how much work is done by the force
ii)the final kinetic energy of the block
iii)the final velocity
iv)Using newton's second law,find the acceleration and then the final velocity
1
Expert's answer
2021-10-18T10:57:59-0400

i)

"W=Fd=12\cdot3=36" J

F - horizontal force, d - distance


ii)

kinetic energy:

"E=W=36" J


iii)

"E=mv^2\/2"

"v=\sqrt{2E/m}=\sqrt{2\cdot36/200}=0.6" m/s


iv)

"ma=F"

"a=F\/m=12\/200=0.06" m/s2

"v=at+v_0"

"d=at^2\/2+v_0t+d_0"

if start velocity "v_0=0" and start distance "d_0=0" , then:

"t=\sqrt{2d/a}"

final velocity:

"v=a\sqrt{2d/a}=0.06\sqrt{2\cdot3/0.06}=0.6" m/s